OpenStudy (anonymous):

How you find the LCM of (4+3x), (16-9x^2), and (4-3x)

OpenStudy (saifoo.khan):

(16-9x^2)

OpenStudy (anonymous):

HI can you show me how to do it?

OpenStudy (saifoo.khan):

LCM is Like whatever we find maximum terms among all three.

OpenStudy (anonymous):

ok sound simple enough

OpenStudy (saifoo.khan):

That's my self invented thing. LOL.

OpenStudy (anonymous):

ahhhhhh...

OpenStudy (anonymous):

For real I would like to know what is the process to get to that.

OpenStudy (saifoo.khan):

there is no particular way to solve that... (4+3x), (16-9x^2), and (4-3x) (4+3x), (4+3x)(4-3x) and (4-3x) Now take out commons..

OpenStudy (saifoo.khan):

common, (4+3x)(4-3x) Finally, (16 - 9x^2)

OpenStudy (cwrw238):

(16-9x^2) = (4-3x)*(4+3x) so the lowest common multiple they will all divide into is (16-9x^2)

OpenStudy (anonymous):

I think I understand

OpenStudy (saifoo.khan):

Cool.. =D please hit good answer.

OpenStudy (cwrw238):

another example is lcm of x, x^2 and x^3 = x^3